OpenStudy (anonymous):

A flower bed has the shape of a rectangle 18 feet long and 12 feet wide. What is its area in square yards?

Vocaloid (vocaloid):

well, first convert 18 feet and 12 feet to yards 1 yard = 3 feet so 18 feet = __ yards 12 feet = ___ yards can you try filling in the blanks?

OpenStudy (anonymous):

6 & 2

OpenStudy (anonymous):

@Vocaloid

Vocaloid (vocaloid):

6 is right 12 feet = __ yards

OpenStudy (anonymous):

4

OpenStudy (anonymous):

@Vocaloid

OpenStudy (tkhunny):

\((18\;ft)*(12\;ft) = 216\;ft^{2}\cdot\left(\dfrac{1\;yd}{3\;ft}\right)^{2}\) Alternatively \(\left(18\;ft\cdot\dfrac{1\;yd}{3\;ft}\right)*\left(12\;ft\cdot\dfrac{1\;yd}{3\;ft}\right) = (6\;yd)*(4\;yd)\) The notation can help you if you let it by practicing.

Vocaloid (vocaloid):

right, so we have 6 yards and 4 yards as our length and width to find the area, all we need to do is multiply 6*4 = ?

OpenStudy (anonymous):

24 @Vocaloid

Vocaloid (vocaloid):

yup so our final answer is 24 yrds^2

OpenStudy (anonymous):

Thank you guys both